Add 15 to a number, then multiply the results by 3. The answer is the same as subtracting 4 from from the number and multiplying the results by 5. Find the number?

Answer by ikleyn(53763)

.

Let x be the (unknown) number.


First sentence is translated into this expression  3*(x+15).


Second sentence is translated into this expression  5*(x-4).


These numbers are the same, so we write equation

    3*(x+15) = 5*(x-4).


Then we solve equation using standard procedures

    3x + 45 = 5x - 20

    45 + 20 = 5x - 3x

      65    =    2x

       x    =    65/2 = 32.5.


ANSWER.  The number is 32.5.


CHECK.  You can easily check that 3(x+5) = 3*(32.5+5) = 142.5 is the same 

        as 3*(x-4) = 3*(32.5-4) = 142.5.    It confirms the answer.

Solved.

---------------------

If you want to learn on how to solve simple and simplest linear equations,  read the lesson
    - HOW TO solve a linear equation
in this site.

Consider this lesson as your textbook,  handbook,  tutorials and  (free of charge)  home teacher.

Happy learning  (!)